Solve the following problem. Be sure to show all steps (V.E.S.T.) and work in order to receive full credit In a triangle, the middle length side is 3 more than the shortest side, and the longest side is double the middle length side. The perimeter of the triangle is 45 cm.
Find the lengths of the sides

Answer by richard1234(7193)

Let x be the shortest length. Then x+3 is the middle length, and 2(x+3) = 2x+6 is the longest side. Then,


x + (x+3) + (2x+6) = 45

4x + 9 = 45

4x = 36, x = 9.

The sides of the triangle are x, x+3, 2x+6, i.e. 9,12,24

Wow, terrible problem. A triangle cannot have lengths of sides 9,12,24 (this violates the triangle inequality). Tell your teacher or professor that the problem epically fails.